# DiffNorm: Self-Supervised Normalization for Non-autoregressive Speech-to-speech Translation

### v1

@@ 1 Introduction
In this work, we propose DiffNorm, a self-supervised speech normalization strategy that alleviates the multi-modality problem of NAT models by simplifying the target distribution. Instead of distilling training data from an autoregressive model [9] or utilizing perturbed speech to train a normalizer [31, 20], we rely on the denoising objective of Denoising Diffusion Probabilistic Models [14, DDPM] to normalize target speech units. DiffNorm inject synthetic noise to speech features and use diffusion model to gradually recover the feature, obtaining a simplified and more consistent data distribution that obscures non-crucial details. As the denoising objective is learned in a self-supervised manner over latent speech representations, it eliminates the necessity for transcription data [31] or manually crafted perturbation functions [20]. As illustrated in Fig. 1, applying DiffNorm to the target data obtains normalized speech units that lead to better NAT training for speech-to-unit translation.
Besides using DiffNorm as a data-centric strategy to mitigate the multi-modality problem, we also propose a regularization strategy to enhance the NAT model’s robustness and generalizability when facing complex data distribution (e.g., linguistic diversity and acoustic variation [20, 19]). Inspired by classifier-free guidance [16, CG], during training, we occasionally drop out source information and replace it with a "null" representation, compelling the models to generate coherent units without conditioning on the source data. During iterative parallel decoding of the NAT model, we obtain higher-quality translation by mixing conditional and unconditional generation. Ultimately, combining DiffNorm and CG results in our top-performing state-of-the-art system, achieving approximately $+7$ and $+2$ ASR-BLEU increment for En-Es and En-Fr translation compared to previous non-autoregressive systems on the CVSS [23] benchmark.
@@ 2 Problem formulation and overview
We aim to develop a direct (textless) speech-to-speech translation system that transduces a source speech $\bm{x}=(x_{1},\cdots,x_{N})$ into target speech. We follow Lee et al., 2022a [30] to reduce speech-to-speech translation into two sub-tasks: speech-to-unit translation and unit-to-speech synthesis. In this work, we focus on speech-to-unit translation and follow prior work [30, 31, 20] to use the same unit-to-speech component for a fair comparison. To generate speech units for the target language, we first extract speech feature $\bm{h}=(h_{1},\cdots,h_{M})\in\mathbb{R}^{M\times H}$, where each feature has $H$ dimensions. Subsequently, a K-means clustering model is trained on extracted features and used to generate speech units $\bm{y}=(y_{1},\cdots,y_{M})\in\mathbb{R}^{M\times 1}$. Once the source speech $\bm{x}$ and target speech unit $\bm{y}$ are prepared, we train sequence-to-sequence models to translate from source speech into target units. In practice, we follow Lee et al., 2022b [31] to use a multilingual-HuBERT (mHuBERT) model to extract $H=768$ dimensional speech features $\bm{h}$. Then we use a 1000-cluster K-means model to predict speech units given the encoded features.222 We take the mHuBERT and K-means models off-the-shelf to produce units consistent with prior work. See github.com/facebookresearch/fairseq/tree/main/examples/speech_to_speech For speech-to-unit translation, we follow Huang et al., [20] to adopt Conditional Masked Language Modeling [9, CMLM], a kind of non-autoregressive transformer (NAT).
@@
Additionally, we propose to incorporate classifier-free guidance [16, CG], a widely adopted strategy for diffusion-based image generation, as a regularization method to improve the non-autoregressive speech-to-unit system (section § 4). As shown in Fig. 3, by forcing NAT models to unmask target speech units without conditioning on source information, the model becomes more robust, generating coherent speech units that result in higher-quality translations. During inference, we follow classifier-free guidance to mix conditional and unconditional generation for the NAT model’s iterative decoding, further enhancing its translation quality.
@@ 3 DiffNorm: denoising diffusion models for speech normalization
Previous normalization strategy [31, 20] on speech-to-unit translation relies on connectionist temporal classification (CTC) fine-tuning [3] using the HuBERT [18] model. For example, Lee et al., 2022b [31] rely on single-speaker data from VoxPopuli [53] to produce normalized (speaker-invariant) speech units for HuBERT-CTC fine-tuning. On the other hand, Huang et al., [20] generate acoustic-agnostic units by perturbing rhythm, pitch, and energy information with (manually) pre-defined transformation functions. We propose to normalize speech units using self-supervised denoising objectives from Denoising Diffusion Probabilistic Models [14, DDPM]. We believe DDPM is more suitable for speech normalization because: (1) It only requires monolingual speech data, without the need of transcriptions to create a text-to-unit model as in [31]. (2) It learns to denoise features in a high-dimensional space rather than relying on hand-designed perturbation as in [20]. DiffNorm consists of a variational auto-encoder (VAE) and a diffusion model. Since vanilla VAE [26] and DDPM [14] are applied to image generation, we modify them to support token generation and incorporate multitasking objectives. The VAE model reduces the dimension of the speech feature, mapping feature $\bm{h}$ into lower dimensional latent $\bm{z}$. The diffusion model (visualized in Fig. 2) is then trained to denoise on the latent representation space, aiming to recover the original latent $\bm{z}_{0}$ from the standard Gaussian noise $\bm{z}_{T}$.333 The subscript $z_{0}$ is added to indicate the time step for diffusion process and $\bm{z}_{0}$ and $\bm{z}$ are equivalent. In the subsequent sections, we begin by outlining the architecture of our VAE model (§ 3.1). Then, we delve into how the diffusion model is trained using the latent variables encoded by the VAE (§ 3.2).
@@ 3.2 Diffusion model for denoising latent speech representation
In our analysis (§ 6.2), we show that adding NLL and reconstruction loss is indeed helpful in improving the diffusion model’s reconstruction quality. Lastly, to parameterize our diffusion model for noise estimation, we modify Diffusion Transformer [36, DiT] to suit our task. For details of our architecture and hyper-parameters, we refer readers to Appendix B.
Inference    For speech normalization, we choose a start time $T\in[0,200]$ that decides the amount of noise to inject for the diffusion model to recover. Given the start time $T$, we follow Denoising Diffusion Implicit Models [46, DDIM] sampler to reverse noised latent $\bm{z}_{T}$ back to $\bm{z}_{0}$. Then, our VAE model converts $\bm{z}_{0}$ back to speech units with its decoder and language modeling head. Our inference procedure is summarized in Alg. 2 and visualized in Fig. 2.
@@ 4 Classifier-free guidance for non-autoregressive transformer
Training    In our CMLM training with classifier-free guidance, we largely adhere to standard CMLM training [20] but introduce a small dropout probability $p_{\text{drop}}=0.15$ for the encoder representations. Specifically, we parameterize the encoder and decoder of the CMLM as $\phi_{\rm enc}$ and $\phi_{\rm dec}$, respectively. The encoder processes the source speech into a representation $\bm{g}=f(\bm{x};\phi_{\rm enc})$. The decoder then predicts the vocabulary distribution $\bm{v}=f(\hat{\bm{y}}|\bm{g};\phi_{\rm dec})$ from $\bm{g}$ and the randomly masked target speech units $\hat{\bm{y}}$. We calculate the NLL loss using $\bm{v}$ and the ground truth units $\bm{y}$ at the masked positions to train the CMLM model. If dropout is applied, the decoder receives a "null" representation $\bm{g}_{\emptyset}$—a randomly-initialized learnable vector—forcing it to rely solely on the target information to unmask units. For more details of CMLM, we refer readers to Appendix C and previous paper [9, 20].
@@ 5.2 Results
abEn-Es En-Fr Speed Speedup
@@
1 Transformer [30] 10.07 15.28 870 1.00$\times$
2 Norm Transformer [31] 12.98 15.93 870 1.00$\times$
3 Conformer 13.75 17.07 895 1.02$\times$
@@
4 CMLM 12.58 15.62 4651 5.34$\times$
5 CMLM + BiP[20] 12.62 16.97
@@
6 CMLM + DiffNorm 18.96 17.27 4651 5.34$\times$
@@
Table 2: Comparison of speech-to-speech models evaluated by quality (ASR-BLEU) and speed (units/seconds). Results with are taken from the prior work [20]. We use $w=0.5$ for CG. Our NAT models achieve superior translation quality while maintaining their fast inference speed.
@@ 6.1 Effect of synthetic noise injection
Start Step Scheduler Parameters En-Es En-Fr
original units N/A N/A N/A 100 48.7 12.58 100 40.64 15.62
$T=50$ 0.007 0.917 0.398 89.4 48.5 18.56 91.1 40.38 17.12
$T=100$ 0.016 0.697 0.717 81.2 47.63 18.96 83.0 39.9 16.69
$T=120$ 0.022 0.577 0.816 75.3 46.25 16.7 77.7 39.03 17.27
$T=150$ 0.038 0.373 0.928 57.8 31.43 14.77 60.0 28.3 14.03
6.2 Ablation on training objectives for DiffNorm
DiffNorm requires an auto-encoder to map speech features into lower-dimension latents and then train denoising diffusion models using the encoded latents. In this section, we perform ablation experiments to investigate the effect of (1) latent dimension (2) Gaussian constraints (3) multitask objective for diffusion models. Firstly, to investigate the effect of latent dimension, we train auto-encoders that encode speech features into 16, 32, and 128 dimensions. As shown in Table 3, it comes as no surprise that a larger latent dimension yields superior reconstruction results, evidenced by higher accuracy across systems trained with varying objectives. Next, we turn our attention to the importance of applying Gaussian constraints to the auto-encoder’s latent representation space. Our results (comparing with and without KL) reveal that incorporating Gaussian constraints is crucial. Latent spaces not regularized by $\mathcal{L}_{\text{kl}}$ lead to significantly poorer performance. Finally, we explore the effectiveness of our proposed multitasking objective for the diffusion model by training another vanilla DDPM solely with $\mathcal{L}_{\rm noise}$ (no Multitask). We find that the diffusion model’s reconstruction capability indeed improves when employing the multitasking objective, particularly when denoising from a more noisy latent representation (e.g., $T=150$). Through our ablation analysis, we identify the optimal setup for our DiffNorm model, which involves (1) mapping speech features to 128 dimensions, (2) regularizing latent space by Gaussian constraints, and (3) utilizing the multitask objective for diffusion modeling training.
@@ 7 Related work
Speech normalization    Inspired by previous work on speech enhancement [49, 1], Lee et al., 2022b [31] propose to normalize speech by synthesizing speaker-invariant waveforms through text-to-speech (TTS) systems [55, 44, 40, 39]. Huang et al., [20] propose to normalize speech with Bilateral Perturbation that focus on the rhythm, pitch, and energy information. Different from previous normalization methods that requires transcription [31] or manually designed perturbation [20], our DiffNorm strategy leverages diffusion models. Diffusion models [14, 46, 47] have achieved remarkable generative ability to produce high-quality images [7, 15, 41, 36, 43] and audio [32, 38, 45]. Using self-supervised denoising objectives [14], we train effective denoising models capable of normalizing speech for training better speech-to-unit systems.
Non-autoregressive speech-to-speech translation    For sequence-to-sequence modeling, auto-regressive [17, 5, 52] and non-autoregressive [11, 12, 9, 19] models have been widely explored. Lee et al., 2022a [30] reduce speech-to-speech into speech-to-unit task and follow the widely-used modeling strategy, Transformer [52], to predict speech units. Later, non-autoregressive models [20, 8] have been explored and Huang et al., [20] is the first to use a non-autoregressive transformer model, Conditional Masked Language Model [9, CMLM], for speech-to-unit task. Fang et al., [8], on the other hand, adopts Directed Acyclic Transformer [19] for speech-to-unit translation.
@@ C.1 Background on Non-Autoregressive Transformers
Non-autoregressive Transformers [11, 12, 9, inter alia.] is a family of models that transduce sequences. In this work, we follow the formalization from the widely-used NAT: Conditional Masked Language Model [9, CMLM]. CMLM consists of an encoder $\phi_{\rm enc}$ that represents the input as high-dimensional features and a decoder $\phi_{\rm dec}$ that generates tokens conditioning on the encoded features. Different from auto-regressive Transformers [52], CMLM’s decoder does not apply causal masking and is trained with an unmasking objective.

### v2

@@ 1 Introduction
In this work, we propose DiffNorm, a self-supervised speech normalization strategy that alleviates the multi-modality problem of NAT models by simplifying the target distribution. Instead of distilling training data from an autoregressive model [10] or utilizing perturbed speech to train a normalizer [32, 21], we rely on the denoising objective of Denoising Diffusion Probabilistic Models [15, DDPM] to normalize target speech units. DiffNorm inject synthetic noise to speech features and use diffusion model to gradually recover the feature, obtaining a simplified and more consistent data distribution that obscures non-crucial details. As the denoising objective is learned in a self-supervised manner over latent speech representations, it eliminates the necessity for transcription data [32] or manually crafted perturbation functions [21]. As illustrated in Fig. 1, applying DiffNorm to the target data obtains normalized speech units that lead to better NAT training for speech-to-unit translation.
Besides using DiffNorm as a data-centric strategy to mitigate the multi-modality problem, we also propose a regularization strategy to enhance the NAT model’s robustness and generalizability when facing complex data distribution (e.g., linguistic diversity and acoustic variation [21, 20]). Inspired by classifier-free guidance [17, CG], during training, we occasionally drop out source information and replace it with a "null" representation, compelling the models to generate coherent units without conditioning on the source data. During iterative parallel decoding of the NAT model, we obtain higher-quality translation by mixing conditional and unconditional generation. Ultimately, combining DiffNorm and CG results in our top-performing state-of-the-art system, achieving approximately $+7$ and $+2$ ASR-BLEU increment for En-Es and En-Fr translation compared to previous non-autoregressive systems on the CVSS [24] benchmark.
@@ 2 Problem formulation and overview
We aim to develop a direct (textless) speech-to-speech translation system that transduces a source speech $\bm{x}=(x_{1},\cdots,x_{N})$ into target speech. We follow Lee et al., 2022a [31] to reduce speech-to-speech translation into two sub-tasks: speech-to-unit translation and unit-to-speech synthesis. In this work, we focus on speech-to-unit translation and follow prior work [31, 32, 21] to use the same unit-to-speech component for a fair comparison. To generate speech units for the target language, we first extract speech feature $\bm{h}=(h_{1},\cdots,h_{M})\in\mathbb{R}^{M\times H}$, where each feature has $H$ dimensions. Subsequently, a K-means clustering model is trained on extracted features and used to generate speech units $\bm{y}=(y_{1},\cdots,y_{M})\in\mathbb{R}^{M\times 1}$. Once the source speech $\bm{x}$ and target speech unit $\bm{y}$ are prepared, we train sequence-to-sequence models to translate from source speech into target units. In practice, we follow Lee et al., 2022b [32] to use a multilingual-HuBERT (mHuBERT) model to extract $H=768$ dimensional speech features $\bm{h}$. Then we use a 1000-cluster K-means model to predict speech units given the encoded features.222 We take the mHuBERT and K-means models off-the-shelf to produce units consistent with prior work. See github.com/facebookresearch/fairseq/tree/main/examples/speech_to_speech For speech-to-unit translation, we follow Huang et al., [21] to adopt Conditional Masked Language Modeling [10, CMLM], a kind of non-autoregressive transformer (NAT).
@@
Additionally, we propose to incorporate classifier-free guidance [17, CG], a widely adopted strategy for diffusion-based image generation, as a regularization method to improve the non-autoregressive speech-to-unit system (section § 4). As shown in Fig. 3, by forcing NAT models to unmask target speech units without conditioning on source information, the model becomes more robust, generating coherent speech units that result in higher-quality translations. During inference, we follow classifier-free guidance to mix conditional and unconditional generation for the NAT model’s iterative decoding, further enhancing its translation quality.
@@ 3 DiffNorm: denoising diffusion models for speech normalization
Previous normalization strategy [32, 21] on speech-to-unit translation relies on connectionist temporal classification (CTC) fine-tuning [4] using the HuBERT [19] model. For example, Lee et al., 2022b [32] rely on single-speaker data from VoxPopuli [58] to produce normalized (speaker-invariant) speech units for HuBERT-CTC fine-tuning. On the other hand, Huang et al., [21] generate acoustic-agnostic units by perturbing rhythm, pitch, and energy information with (manually) pre-defined transformation functions. We propose to normalize speech units using self-supervised denoising objectives from Denoising Diffusion Probabilistic Models [15, DDPM]. We believe DDPM is more suitable for speech normalization because: (1) It only requires monolingual speech data, without the need of transcriptions to create a text-to-unit model as in [32]. (2) It learns to denoise features in a high-dimensional space rather than relying on hand-designed perturbation as in [21]. DiffNorm consists of a variational auto-encoder (VAE) and a diffusion model. Since vanilla VAE [27] and DDPM [15] are applied to image generation, we modify them to support token generation and incorporate multitasking objectives. The VAE model reduces the dimension of the speech feature, mapping feature $\bm{h}$ into lower dimensional latent $\bm{z}$. The diffusion model (visualized in Fig. 2) is then trained to denoise on the latent representation space, aiming to recover the original latent $\bm{z}_{0}$ from the standard Gaussian noise $\bm{z}_{T}$.333 The subscript $z_{0}$ is added to indicate the time step for diffusion process and $\bm{z}_{0}$ and $\bm{z}$ are equivalent. In the subsequent sections, we begin by outlining the architecture of our VAE model (§ 3.1). Then, we delve into how the diffusion model is trained using the latent variables encoded by the VAE (§ 3.2).
@@ 3.2 Diffusion model for denoising latent speech representation
In our analysis (§ 6.2), we show that adding NLL and reconstruction loss is indeed helpful in improving the diffusion model’s reconstruction quality. Lastly, to parameterize our diffusion model for noise estimation, we modify Diffusion Transformer [37, DiT] to suit our task. For details of our architecture and hyper-parameters, we refer readers to Appendix B.
Inference    For speech normalization, we choose a start time $T\in[0,200]$ that decides the amount of noise to inject for the diffusion model to recover. Given the start time $T$, we follow Denoising Diffusion Implicit Models [48, DDIM] sampler to reverse noised latent $\bm{z}_{T}$ back to $\bm{z}_{0}$. Then, our VAE model converts $\bm{z}_{0}$ back to speech units with its decoder and language modeling head. Our inference procedure is summarized in Alg. 2 and visualized in Fig. 2.
@@ 4 Classifier-free guidance for non-autoregressive transformer
Training    We largely adhere to standard CMLM training [21] but introduce a small dropout probability $p_{\text{drop}}=0.15$ for the encoder representations. Specifically, we parameterize the encoder and decoder of the CMLM as $\phi_{\rm enc}$ and $\phi_{\rm dec}$, respectively. The encoder processes the source speech into a representation $\bm{g}=f(\bm{x};\phi_{\rm enc})$. The decoder then predicts the vocabulary distribution $\bm{v}=f(\hat{\bm{y}}|\bm{g};\phi_{\rm dec})$ from $\bm{g}$ and the randomly masked target speech units $\hat{\bm{y}}$. The total amount of masked token is uniformed sampled: $n\sim\mathcal{U}[1,M]$; subsequently, $n$ of $M$ tokens from the target units $\bm{y}$ are randomly masked to form noisy target units $\hat{\bm{y}}$. CMLM also trains a length predictor that estimates the output length given input sequences and we refer readers to [21] for more details.
Using the predicted distribution $\bm{v}$, we compute the NLL loss against the ground truth units $\bm{y}$ at the masked positions to train the model. If dropout is applied, the decoder receives a "null" representation $\bm{g}_{\emptyset}$—a randomly-initialized learnable vector—forcing it to rely solely on the target information to unmask units. For more details, we refer readers to Appendix C and previous paper [10, 21].
@@ 5.2 Results
abEn-Es En-Fr Fr-En Speed Speedup
@@
1 Transformer [31] 10.07 15.28 15.44 870 1.00$\times$
2 Norm Transformer [32] 12.98 15.93 15.81 870 1.00$\times$
3 Conformer 13.75 17.07 18.02 895 1.02$\times$
@@
4 CMLM 12.58 15.62 16.95 4651 5.34$\times$
5 CMLM + BiP[21] 12.62 16.97 18.03
@@
6 CMLM + DiffNorm 18.96 17.27 19.53 4651 5.34$\times$
@@
Table 2: Comparison of speech-to-speech models evaluated by quality (ASR-BLEU) and speed (units/seconds).: Fr-En experiments are added during author response period and we leave model 7,8 for future work. Results with are taken from the prior work [21]. We use $w=0.5$ for CG. Our NAT models achieve superior translation quality while maintaining their fast inference speed.
@@ 6.1 Effect of synthetic noise injection
Start Step Scheduler Parameters En-Es En-Fr
original units N/A N/A N/A 100 48.7 12.58 100 40.64 15.62
$T=50$ 0.007 0.917 0.398 89.4 48.5 18.56 91.1 40.38 17.12
$T=100$ 0.016 0.697 0.717 81.2 47.63 18.96 83.0 39.9 16.69
$T=120$ 0.022 0.577 0.816 75.3 46.25 16.7 77.7 39.03 17.27
$T=150$ 0.038 0.373 0.928 57.8 31.43 14.77 60.0 28.3 14.03
Start Step Acc-Rec BL-Rec
$T=10$ 93.8 15.98
$T=30$ 91.6 17.92
Table 4: Reconstruction and downstream performance with small noise injection.
From the Table 3, we observed a substantial enhancement in En-Es translation performance with $T=50$. Consequently, we conducted additional experiments for En-Es translation with start times $T=10$ and $T=30$, detailed in Table 4. The results show that at a minimal start time of $T=10$, the reconstruction accuracy reaches 93.8%, yielding a downstream ASR-BLEU score of 15.98. This score significantly surpasses the baseline CMLM result of 12.58. The high accuracy indicates that about 6% of tokens vary post-reconstruction, likely due to the VAE model’s regularization effect since the diffusion model does not cause notable reconstruction deviations at such a small $T$. This suggests that the Spanish speech dataset might have considerable acoustic variations and noise, which the VAE model can partially mitigate. As $T$ increases, the diffusion model further refines the representation, leading to improved downstream results, as evidenced by the rise in ASR-BLEU score from from $T=10$ to $T=100$. This ablation study underscores that the optimal choice of T is affected by the dataset quality.
6.2 Ablation on training objectives for DiffNorm
@@
DiffNorm requires an auto-encoder to map speech features into lower-dimension latents and then train denoising diffusion models using the encoded latents. In this section, we perform ablation experiments to investigate the effect of (1) latent dimension (2) Gaussian constraints (3) multitask objective for diffusion models. Firstly, to investigate the effect of latent dimension, we train auto-encoders that encode speech features into 16, 32, and 128 dimensions. As shown in Table 5, it comes as no surprise that a larger latent dimension yields superior reconstruction results, evidenced by higher accuracy across systems trained with varying objectives. Next, we turn our attention to the importance of applying Gaussian constraints to the auto-encoder’s latent representation space. Our results (comparing with and without KL) reveal that incorporating Gaussian constraints is crucial. Latent spaces not regularized by $\mathcal{L}_{\text{kl}}$ lead to significantly poorer performance. Finally, we explore the effectiveness of our proposed multitasking objective for the diffusion model by training another vanilla DDPM solely with $\mathcal{L}_{\rm noise}$ (no Multitask). We find that the diffusion model’s reconstruction capability indeed improves when employing the multitasking objective, particularly when denoising from a more noisy latent representation (e.g., $T=150$). Through our ablation analysis, we identify the optimal setup for our DiffNorm model, which involves (1) mapping speech features to 128 dimensions, (2) regularizing latent space by Gaussian constraints, and (3) utilizing the multitask objective for diffusion training.
@@ 7 Related work
Speech normalization    Speech representation are typically extracted from pre-trained encoders [19, 3, 40, 5] and can be compressed or adapted in different speech-to-speech/text tasks [61, 63, 50, 52, 51]. Inspired by previous work on speech enhancement [54, 1], Lee et al., 2022b [32] propose to normalize speech by synthesizing speaker-invariant waveforms through text-to-speech (TTS) systems [60, 46, 42, 41]. Huang et al., [21] propose to normalize speech with Bilateral Perturbation that focus on the rhythm, pitch, and energy information. Different from previous normalization methods that requires transcription [32] or manually designed perturbation [21], our DiffNorm strategy leverages diffusion models. Diffusion models [15, 48, 49] have achieved remarkable generative ability to produce high-quality images [8, 16, 43, 37, 45] and audio [33, 39, 47]. Using self-supervised denoising objectives [15], we train effective denoising models capable of normalizing speech for training better speech-to-unit systems.
Non-autoregressive speech-to-speech translation    For sequence-to-sequence modeling, auto-regressive [18, 6, 57] and non-autoregressive [12, 13, 10, 20] models have been widely explored. Lee et al., 2022a [31] reduce speech-to-speech into speech-to-unit task and follow the widely-used modeling strategy, Transformer [57], to predict speech units. Later, non-autoregressive models [21, 9] have been explored and Huang et al., [21] is the first to use a non-autoregressive transformer model, Conditional Masked Language Model [10, CMLM], for speech-to-unit task. Fang et al., [9], on the other hand, adopts Directed Acyclic Transformer [20] for speech-to-unit translation.
@@ 8 Conclusion
Acknowledgement
We express our profound appreciation to anonymous reviewers for their helpful suggestions. We also thank Xiuyu Li for his valuable suggestions, which greatly enriched our work. Lastly, we thank JHU + Amazon Initiative for Interactive AI for sponsoring the work.
@@ C.1 Background on Non-Autoregressive Transformers
Non-autoregressive Transformers [12, 13, 10, inter alia.] is a family of models that transduce sequences. In this work, we follow the formalization from the widely-used NAT: Conditional Masked Language Model [10, CMLM]. CMLM consists of an encoder $\phi_{\rm enc}$ that represents the input as high-dimensional features and a decoder $\phi_{\rm dec}$ that generates tokens conditioning on the encoded features. Different from auto-regressive Transformers [57], CMLM’s decoder does not apply causal masking and is trained with an unmasking objective.